OpenStudy (anonymous):

Mrs. Ferber graphed a line in class for her very bright students. The line passed through the points (6,-1) and (-2,3). She asked them to find the equation for the line and write it in standard form, Ax+By=C, with A≥0. Her brilliant students all correctly wrote the equation. What was the value of B?

OpenStudy (texaschic101):

first we find the slope slope(m) = (y2 - y1) / (x2 - x1) (6,-1) x1 = 6 and y1 = -1 (-2,3) x2 = -2 and y2 = 3 now sub slope(m) = (3 - (-1) / (-2 - 6) slope(m) = (3 + 1) / -8 slope(m) = 4/-8 or -4/8 which reduces to -1/2 so our slope is -1/2 now we will use y = mx + b slope(m) = -1/2 you can use either of your sets of points...I will use (6,-1) x = 6 and y = -1 now sub and solve for b -1 = -1/2(6) + b -1 = -3 + b -1 + 3 = b 2 = b so our equation is : y = -1/2x + 2, but we want it in standard form Ax + By = C y = -1/2x + 2 --- multiply the entire equation by 2 2y = -x + 4 -- add x to both sides x + 2y = 4 <=== standard form So the value of B is 2

OpenStudy (anonymous):

wow thanks you!!!!

OpenStudy (texaschic101):

sure thing :)

OpenStudy (lena772):

Welcome to OpenStudy. Please take some time to see our TaC and CoC. http://openstudy.com/terms-and-conditions http://openstudy.com/code-of-conduct